# A New Dataset for Monocular Depth Estimation Under Viewpoint Shifts

### v1

@@ Abstract
Keywords:
Dataset Domain Shift Monocular Depth Estimation
1 Introduction
@@
Recent years have seen significant advancements in the machine learning community, with transformers in particular achieving state-of-the-art results on multiple tasks [22, 6]. Autonomous driving (AD) has been receiving more and more attention from the public with different forms of it being regulated and approved for public implementations, from fully autonomous robotaxis [3] to consumer vehicles with different levels of autonomy [1, 2]. Perception, and computer vision in particular, is the most important building block for these technologies. A fundamental task for the perception of the surroundings is depth estimation. It is well known that such tasks are susceptible to out-of-distribution samples, from adverse weather and light conditions to image corruptions [19, 24, 16].
@@
Another distribution shift that is common in this context is viewpoint change, either caused by vehicles with different sizes or by variations in the camera installation. A few works have recently started to shed light on the effect of such shifts on related tasks like birds-eye-view semantic segmentation and object detection [15, 23]. However, quantification of such an effect on monocular depth estimators remains heavily underexplored, with only [21] giving some insights on the behaviours of these models. This could be mainly due to the lack of annotated data and the cost and technical difficulties associated with it. Traditionally, this kind of task is performed using measurements from expensive lidar sensors [8, 26, 5]. These sensors are not ubiquitous and require multiple layers of post-processing, such as outlier detection and removal, data alignment, registration with other sources, and other computationally expensive steps. A possible solution lies in the domain of synthetic data, which introduces a well-known sim-to-real domain shift [12].
@@
Our proposal is to remove the dependency on expensive sensors to evaluate depth estimation techniques by exploiting basic geometric elements such as homographies. We estimate a ground truth (GT) homography from an initial calibration session that maps points from the image plane into the road plane. We use the homography to compute distances of road objects that serve as GT for our evaluation. As a matter of fact, our proposal is to use object distance estimation as a proxy task for traditional depth evaluation. We are aware that this approach is accurate only if the geometry of the road plane is in line with the one measured during calibration, otherwise introducing an intrinsic error. However, we empirically show that such an error is considerably limited on planar geographies such as the one from the KITTI dataset [8]. Furthermore, experiments on KITTI show that our proxy task provides evaluations that are very close to the ones observed using lidar GT.
@@
Our strategy enabled us to collect a large dataset of video sequences, recorded from different known and calibrated camera settings. We use such data to highlight the impact of the different viewpoint shifts on a modern depth estimator and give insights on their effects on the scale perceived by the model. The dataset can be accessed at kaggle.com/datasets/aurelpjetri/viewpointdepth.
@@
  • we discuss in-depth the limitations of the proposed approach and empirically show its effectiveness on a public annotated dataset;
  • we release a new dataset of road scenes collected from multiple viewpoints;
@@
The remainder of this paper is structured as follows: Sec. 2 presents the most relevant works in this context, Sec. 3 outlines the proposed ground truth and evaluation strategy, Sec. 4 presents the details of the collected dataset, Sec. 5 describes the different experiments conducted and Sec. 6 draws final conclusions.
2 Related Work
In this section we review work from the recent literature most related to our contributions.
2.1 Road Scene Depth Datasets
@@
2.2 Self-supervised Monocular Depth Estimation
Depth estimation is a fundamental geometric task for 3D scene understanding. Self-supervised monocular depth estimation is a particular variant of this task that enables training networks without the use of explicit depth labels. Although it is an ill-posed problem, since multiple 3D scenes can generate the same 2D image, some deep learning techniques have shown good results. One of the first proposed approaches used the pretext task of reconstructing frames in a video sequence to force the network to learn a good depth estimation of the scene [29]. The depth network is trained alongside a pose network for the new frame to be rendered. Given a source frame at time $t$ and a target frame at $t+1$, each pixel $p_{t+1}$ can be back-projected to the source view using depth $D_{t}$ and pose $P_{t+1\rightarrow t}$ predicted by the networks:
$p_{t}\sim KP_{t+1\rightarrow t}K^{-1}p_{t+1}.$ (1)
In this way bilinear sampling can be used to synthesise $\hat{p}_{t+1}$ [13]. A direct comparison of each pixel $i$ within each frame $t$ for training loss can then be used [29]:
$\mathcal{L}=\sum_{t}\sum_{i}\left|p^{i}_{t}-\hat{p}^{i}_{t}\right|.$ (2)
Inspired by this work, [10] introduced multiple improvements, such as the use of a photometric error in the loss function combining L1 and SSIM [25]. The results were impressive, not only because they made state-of-the-art, but also given how close they were to stereo-based methods, proving the effectiveness of this paradigm.
@@ 2.3 Viewpoint Shift
3 Ground Truth and Evaluation Strategy
@@ 3.1 Calibration
The calibration phase is the only manual step of the method. We select reference ground points and measure their distance to the camera to construct the target plane with a metric reference system centered in the camera (see Fig. 1), so that for any point $\text{X}=[x,y]^{T}$ its distance to the camera can be computed by simply $\sqrt{x^{2}+y^{2}+h^{2}}$, where $h$ is the height of the camera. This allows us to estimate a homography mapping points in the image to points on the target plane, and therefore to associate a metric distance to any point in the image, provided that it lies on the ground. Given a point on the image $\text{x}=[u,v]^{T}$, we project it to X in the target plane:
$[X,Y,W]^{T}=H\cdot[u,v,1]^{T}\;\Rightarrow\;\text{X}=[X/W,Y/W]^{T}=[x,y]^{T}.$ (4)
Figure 1: 1(a) Manually labeled points of a parking lot pattern (source points for the homography). 1(b) Target plane with metric distances (target points for the homography).
3.2 Object Detection
@@
We use an object detector to detect the relevant road objects in the image: car, truck, bus, motorcycle, bicycle and person. Given a bounding box with coordinates $(x_{1},y_{1})$ and $(x_{2},y_{2})$, we assume that all objects lie on the ground, and therefore we define their distance as that of the center of the lower side of the box $\hat{X}=((x_{2}-x_{1})/2,y_{2})$ (Fig. 2(a)). By projecting this point using Eq. 4, we are able to associate a metric distance to each object and use it as ground truth for the model evaluation in the next step.
@@
Figure 2: 2(a) Bounding box of a vehicle with the point $\hat{X}$ used for GT distance. 2(b) Depth prediction with bounding box. The original box was resized to $\alpha=75\%$ (dashed line) to extract the percentile $\beta$ of the prediction.
3.3 Evaluation
@@
To measure the performance of a depth estimation model, we only consider road objects. Therefore, we again use the bounding box information from the object detector and use the depth estimation model to infer the dense distance of the points within the boxes. We then select a percentile $\beta$ from the estimated distances and use that as the inferred object distance. Moreover, to account for inaccuracies we resize bounding boxes to a certain fraction $\alpha$ of the original dimension before collecting inferred depths, as shown in Fig. 2(b). $\alpha$ and $\beta$ are treated as hyperparameters that we select in the experimental phase (see Sec. 5.1).
4 Dataset
@@
The aim of this dataset is to measure the effects of different viewpoint shifts on monocular depth estimation models. We collected the data using a 2023 RAM Promaster 2500 cargo van located in Alpharetta (Georgia), United States. Between March 2023 and February 2024, we installed two identical dashcams in the windshield of the vehicle. The first camera, to which we refer as base, was fixed in an ideal position for a dashcam, i.e., close to the rear-view mirror at a height of 177.8 cm with an orientation of 0° roll, 0° yaw and -4° pitch (later moved to 6° pitch as shown in Tab. 1). Example frames recorded from these positions are reported in Fig. 4.
The second camera, which we identify as shifted, was instead used to simulate different viewpoint shifts and was moved on a weekly basis. Tables 1 and 3 show the different combinations of positions and orientations of the cameras during the collection period. For the camera position in the windshield (X, Y, Z) we used the reference system illustrated in Fig. 3(a) centered in the position of the base camera.
Figure 3: 3(a) Reference system on the windshield centered on the base camera, where the Z axis orientation is in the direction of movement of the vehicle. The red dots indicate the different positions of the shifted camera. 3(b) Axes and rotation reference systems following the right-hand rule.
Table 1: Combinations of different positions in the windshield and orientation angles (in degrees). X, Y, and Z represent the position of the camera in the windshield (in centimeters). Objects represents the number of objects detected in the associated frames. For each row we also present the pitch the base camera had during the recording of that viewpoint and he associated objects.
Shifted Dashcam Base Dashcam
ID X Y Z Pitch Yaw Roll Objects Pitch Objects
0 0 30 0 -4 0.0 0 5291 -4 3026
1 5 0 0 -10 0.0 0 4325 6 4146
2 5 0 0 16 0.0 0 5004 6 5458
3 5 0 0 -4 2.5 0 2906 6 2824
4 5 0 0 -4 5.0 0 2281 6 2219
5 5 0 0 16 5.0 0 3068 6 4207
6 -65 -5 -10 5 0.0 -5 4425 6 3916
7 -65 -5 -10 16 0.0 -5 3537 6 3759
8 -65 30 -10 5 0.0 -5 3465 6 3278
9 -65 30 -10 16 0.0 -5 3712 6 4413
It is worth noting that the yaw angle, present in the camera when installed on the left side (lines 6–9 in Tab. 1), was implicitly generated by the curvature of the windshield itself. Roll angles were computed by manually measuring the camera inclination on the windshield. On the other hand, we measured pitch and yaw angles for each viewpoint by manually labeling the horizontal vanishing point from the straight lines on the road $(v_{u},v_{v})$, and computing the angle given the focal length $(f_{x},f_{y})$ of the camera and its principal point $(c_{u},c_{v})$:
pitch $=$ $\frac{180}{\pi}\mbox{arctan}\left(\frac{c_{v}-v_{v}}{f_{y}}\right)$ (5)
yaw $=$ $\frac{180}{\pi}\mbox{arctan}\left(\frac{c_{u}-v_{u}}{f_{x}}\right)$ (6)
@@
A total of 10 different viewpoints were collected. For each viewpoint we release 3 videos of 10 minutes each from different scenes. Given the long time range in which the videos are distributed, we manually selected samples to represent a similar distribution of suburban areas in non-adverse weather and at day-time as shown in Figs. 5 and 4. Dashcams recorded at an average of 30 frames per second in 720p MP4 format and with a 150° diagonal field of view. We down-sampled the data at 10 fps and extracted the frames undistorted in JPEG format, with a quality factor of 92 and 4:2:0 chroma sub-sampling. This resulted in a total of 360K frames.
Figure 4: 4(a) Example of base camera with -4° of pitch. 4(b) Example of base camera with 6° of pitch.
Figure 5: Example frames from the shifted camera positions following the ID in Tab. 1.
Note that although the two cameras are simultaneously recording the same scene, no synchronization mechanism was implemented. This can result in the two recordings being shifted by a variable number of frames. Finally, to foster further research, we also release GPS points and accelerometer data at 1 and 100 Hz, respectively.
4.1 Object Detection
@@
We used a Yolov5 object detector [14], trained on a combination of BDD100K [27] and a private dataset collected with same camera model of this work. We do not expect the object detector to be affected by the viewpoint shifts as both datasets used for training have non-fixed viewpoints. Moreover, from a qualitative analysis, we observe that the combinations in Tab. 1 do not induce a significant change in appearance of the objects. This is also confirmed by smaller-scale experiments that showed no evident change in mean average precision between the different viewpoints.
@@
4.2 Homography
As mentioned in Sec. 3, in order to be able to associate a ground truth distance to road objects, we estimated homographies mapping image ground points to a target plane with metric reference system (Fig. 1) that we release alongside the dataset.
4.2.1 Base Camera Homography
@@
4.2.2 Shifted Camera Homography
@@
5 Experimental Results
In this section we report the results of our experiments. In particular, we validate our method as ground truth source by comparing it with lidar readings on the KITTI dataset in Sec. 5.1. Then, in Sec. 5.2, we use our ground truth method to measure the effects of the different viewpoints on a self-supervised monocular depth estimator.
@@ 5 Experimental Results
5.1 Homography Evaluation on KITTI
The goal of this experiment is to assess the margin of error of our geometry-based strategy by comparing it with lidar ground truth. We take the MonoViT model trained on the KITTI dataset [8] for self-supervised monocular depth estimation and measure its absolute relative error (Eq. 3) using two ground truth sources: laser scans and our strategy.
@@ 5.1.1 Road Geometry
5.2 Viewpoint Shift Effects
@@
As mentioned in Sec. 5.1, predictions need to be scaled to metric distances for evaluation. However, in order to highlight the effects of the shifts on the scale perceived by the model, we choose to use the same scaling factor for all positions. This factor is computed exploiting all the objects $O$ in the training set $T$, both in the zero-shot and fine-tuning experiments:
@@
Table 2 gives the absolute relative errors of the model in the different positions, both in zero-shot and fine-tuned inference. To isolate the effect of the viewpoint, for each position, we test on footage from the same scene recorded by the two cameras simultaneously, and compute the difference in performance between shifted (S) and base (B) camera. The rows with the highest difference between the S and B cameras correspond to the installation positions with the strongest performance degradation.
Table 2: Absolute relative error of MonoViT in different camera positions for zero-shot and fine-tuned networks. We report errors on the shifted (S) and base (B) cameras corresponding to the same scene. Base camera with pitch -4.0°, all other rows are with base camera at pitch 6°. The top 2 largest degradations are in bold, the second 2 are underlined.
@@
Shifted Dashcam Position Zero-shot Fine-tuned
ID X Y Z Pitch Roll Yaw B S S-B B S S-B
0 0 30 0 -4 0 0 18.3 19.4 1.1 16.9 17.3 0.4
1 5 0 0 -10 0 0 19.4 29.0 9.6 18.0 22.9 5.0
2 5 0 0 16 0 0 17.2 26.1 8.8 17.0 21.1 4.1
3 5 0 0 -4 2.5 0 18.5 21.4 2.9 15.1 16.4 1.2
4 5 0 0 -4 5 0 20.9 21.3 0.4 17.2 17.7 0.5
5 5 0 0 16 5 0 20.5 40.4 19.9 19.7 35.8 16.2
6 -65 -5 -10 5 0 -5 19.6 24.0 4.4 16.5 18.1 1.6
7 -65 -5 -10 16 0 -5 19.0 47.5 28.5 17.7 44.1 26.4
8 -65 30 -10 5 0 -5 18.5 23.6 5.1 17.3 21.2 4.0
9 -65 30 -10 16 0 -5 19.2 25.9 6.6 15.1 17.6 2.5
@@
From the zero-shot results, we observe that the two most affected positions are 7 and 5 (in decreasing order). The combination of pitch and yaw has the most degrading effect, leading to more than doubling the error. Pitch and roll have also the effect of doubling the error of the model. Fine-tuning reduces the error in general, however the degrading effects of the 2 worst positions remain almost unchanged. In general we observe that pure rotations or translations don’t affect the model as much as a combination of them, with some more impactful than others.
5.2.1 Scale Effect
@@
Given the single scale factor used for all positions, we believe that the main effect of viewpoint shifts is the distortion of the scale perceived by the model, which could not have been detected if per-image or per-position scaling was used. In order to better investigate this phenomenon, we compute scaling factors as in Eq. 7 on the different positions separately. Tab. 3 reports the computed scales on the zero-shot experiment, along with the error degradation from Tab. 2. A correlation between performance degradation and perceived scale can be seen. In particular, we observe that higher pitch angles correspond to smaller scales perceived by the model (and therefore a bigger scaling factor necessary to align to GT). In the same way, lower pitch angles generate bigger scales perceived. This observation seems to confirm that indeed depth models use vertical position as depth cue for objects as stated by [21]. In general this suggests that adaptation methods tackling this problem should focus on automatic scaling of models at inference time.
Table 3: Scale perceived by the network in different positions. Scale B and Scale S are the scales perceived in the base and shifted cameras. Scale S-B is computed as Scale S - Scale B to highlight the effect of the viewpoint shift. The difference in error (Error S-B) is also reported from Tab. 2 for reference. The top 2 largest degradations are in bold, the second 2 are underlined.
@@
Shifted Dashcam Position Zero-shot
ID X Y Z Pitch Roll Yaw Error S-B Scale B Scale S Scale S-B
0 0 30 0 -4 0.0 0 1.1 3116 3064 -52
1 5 0 0 -10 0.0 0 9.6 3359 2669 -689
2 5 0 0 16 0.0 0 8.8 3167 4313 1145
3 5 0 0 -4 2.5 0 2.9 3310 3086 -224
4 5 0 0 -4 5.0 0 0.4 3580 3327 -252
5 5 0 0 16 5.0 0 19.9 3330 5254 1923
6 -65 -5 -10 5 0.0 -5 4.4 3315 3034 -280
7 -65 -5 -10 16 0.0 -5 28.5 3310 6133 2822
8 -65 30 -10 5 0.0 -5 5.1 3158 2706 -451
9 -65 30 -10 16 0.0 -5 6.6 3397 3970 572
6 Conclusion

### v3

# ViewpointDepth: A New Dataset for Monocular Depth Estimation Under Viewpoint Shifts

@@ Abstract
I Introduction
@@
Autonomous driving has been receiving more and more attention from the public with different forms of it being regulated and approved for public implementations, from fully autonomous robotaxis [1] to consumer vehicles with different levels of autonomy [2, 3]. Perception, and computer vision in particular, is the most important building block for these technologies. A fundamental task for safety and perception of surroundings is depth estimation. Despite the recent advancements in machine learning, it is well known that such tasks are susceptible to out-of-distribution samples, from adverse weather and light conditions to image corruptions [4, 5, 6].
@@
Viewpoint change is another important distribution shift for depth estimation, either caused by vehicles with different sizes or by variations in the camera installation. A few works have only recently started to shed light on the effect of such shifts on related tasks like Bird’s Eye View (BEV) semantic segmentation and object detection [7, 8]. However, quantification of such an effect on monocular depth estimators remains heavily underexplored, with only [9] giving some insights on the behaviours of these models. This could be mainly due to the lack of annotated data and the cost and technical difficulties associated with it. Traditionally, this kind of task is performed using measurements from expensive LIDAR sensors [10, 11, 12]. These sensors are not ubiquitous and require multiple layers of post-processing, such as outlier detection and removal, data alignment, registration with other sources, and other computationally expensive steps. A possible solution lies in the domain of synthetic data, which introduces a well-known sim-to-real domain shift [13].
@@
Our proposal is to remove the dependency on expensive sensors to evaluate depth estimation techniques by exploiting basic geometric elements such as homographies. We estimate a ground truth (GT) homography from an initial calibration session that maps points from the image plane into the road plane (see Fig. 1). Then, we use the homography to compute distances of road objects that serve as GT for our evaluation.
Figure 1: LABEL:sub@fig:hom_a Manually labeled source points for the homography. LABEL:sub@fig:hom_b Target points for the homography with metric distances.
Our proposal is to use object distance estimation as a proxy task for traditional depth evaluation. Although this approach is accurate only if the geometry of the road plane is in line with the one measured during calibration, we empirically show that estimation error is considerably limited on planar road scenes such as those in the KITTI dataset [10]. Furthermore, experiments on KITTI show that our proxy task provides evaluations that are very close those observed using LIDAR GT.
@@
Our strategy enabled us to collect a large dataset of video sequences, recorded from different known and calibrated camera settings. We use such data to highlight the impact of the different viewpoint shifts on a modern depth estimator and to give insights into their effects on the scale perceived by the model.
@@
  • we discuss in depth the limitations of the proposed approach and empirically show its effectiveness on a public annotated dataset;
  • we release a new dataset of road scenes collected from multiple viewpoints; and
@@
II Related Work
II-A Road Scene Depth Datasets
@@
II-B Self-supervised Monocular Depth Estimation
Depth estimation is a fundamental geometric task for 3D scene understanding. Self-supervised monocular depth estimation is a particular variant of this task that enables training networks without the use of explicit depth labels. Although it is an ill-posed problem, since multiple 3D scenes can generate the same 2D image, some deep learning techniques have shown good results. One of the first proposed approaches used the pretext task of reconstructing frames in a video sequence to force the network to learn a good depth estimation of the scene [15]. The depth network is trained alongside a pose network for the new frame to be rendered. Given a source frame at time $t$ and a target frame at $t+1$, each pixel $p_{t+1}$ can be back-projected to the source view using depth $D_{t}$ and pose $P_{t+1\rightarrow t}$ predicted by the networks: $p_{t}\sim KP_{t+1\rightarrow t}K^{-1}p_{t+1}$. In this way bilinear sampling can be used to synthesise $\hat{p}_{t+1}$ [16], and a direct comparison of each pixel within each frame can then be used for training loss [15]. Inspired by this work, [17] introduced multiple improvements, such as the use of a photometric error in the loss function combining L1 and SSIM [18]. The results were impressive, not only because they made state-of-the-art, but also given how close they were to stereo-based methods, proving the effectiveness of this paradigm.
@@ II-C Viewpoint Shift
III Ground Truth and Evaluation Strategy
@@ III-A Calibration
The calibration phase is the only manual step of the method. We select reference ground points and measure their distance to the camera to construct the target plane with a metric reference system centered in the camera (see Fig. 1), so that for any point $\text{X}=[x,y]^{T}$ its distance to the camera can be computed by simply $\sqrt{x^{2}+y^{2}+h^{2}}$, where $h$ is the height of the camera. This allows us to estimate a homography mapping points in the image to points on the target plane, and therefore to associate a metric distance to any point in the image, provided that it lies on the ground.
III-B Object Detection
@@
We use an object detector to detect the relevant road objects in the image: car, truck, bus, motorcycle, bicycle and person. Given a bounding box with coordinates $(x_{1},y_{1})$ and $(x_{2},y_{2})$, we assume that all objects lie on the ground, and therefore we define their distance as that of the center of the lower side of the box $\hat{X}=((x_{2}-x_{1})/2,y_{2})$ (Fig. 2a). By projecting this point we are able to associate a metric distance to each object and use it as ground truth for the model evaluation in the next step.
@@
Figure 2: LABEL:sub@fig:box_a Bounding box of a vehicle with the point $\hat{X}$ used for GT distance. LABEL:sub@fig:box_b Depth prediction. The original box was resized to $\alpha=75\%$ to extract the percentile $\beta$ of the prediction.
III-C Evaluation
@@
To measure the performance of a depth estimation model, we only consider road objects. Therefore, we again use the bounding box information from the object detector and use the depth estimation model to infer the dense distance of the points within the boxes. We then select a percentile $\beta$ from the estimated distances and use that as the inferred object distance. Moreover, to account for inaccuracies we resize bounding boxes to a certain fraction $\alpha$ of the original dimension before collecting inferred depths, as shown in Fig. 2b. $\alpha$ and $\beta$ are treated as hyperparameters that we select in the experimental phase (see Section V-A).
IV Dataset
@@
The aim of this dataset is to measure the effects of different viewpoint shifts on monocular depth estimation models. We collected the data using a 2023 RAM Promaster 2500 cargo van located in Alpharetta (Georgia), United States. Between March 2023 and February 2024, we installed two identical dashcams in the windshield of the vehicle. The first camera, to which we refer as base, was fixed in an ideal position for a dashcam, i.e., close to the rear-view mirror at a height of 177.8 cm with an orientation of 0° roll, 0° yaw and -4° pitch (later moved to 6° pitch as shown in Table I).
The second camera, which we identify as shifted, was instead used to simulate different viewpoint shifts and was moved on a weekly basis. Table I shows the different combinations of positions and orientations of the cameras during the collection period. For the camera position in the windshield (X, Y, Z) we centered the reference system on the base camera on the inside of the car looking forward in the direction of movement, with the X axis oriented horizontally and positive on the right, the Y axis vertically and positive downward and the Z axis oriented in the direction of movement. Moreover roll pitch and yaw are oriented following the right-hand rule. Roll angles were computed by manually measuring the camera inclination on the windshield. On the other hand, we measured pitch and yaw angles for each viewpoint by manually labeling the horizontal vanishing point from the straight lines on the road $(v_{u},v_{v})$, and computing the angle given the focal length $(f_{x},f_{y})$ of the camera and its principal point $(c_{u},c_{v})$, e.g. $(180/\pi)arctan((c_{v}-v_{v})/f_{y})$ for pitch.
[t] Shifted Dashcam Base Dashcam ID X Y Z Pitch Yaw Roll Objects Pitch Objects 0 0 30 0 -4 0.0 0 5291 -4 3026 1 5 0 0 -10 0.0 0 4325 6 4146 2 5 0 0 16 0.0 0 5004 6 5458 3 5 0 0 -4 2.5 0 2906 6 2824 4 5 0 0 -4 5.0 0 2281 6 2219 5 5 0 0 16 5.0 0 3068 6 4207 6 -65 -5 -10 5 0.0 -5 4425 6 3916 7 -65 -5 -10 16 0.0 -5 3537 6 3759 8 -65 30 -10 5 0.0 -5 3465 6 3278 9 -65 30 -10 16 0.0 -5 3712 6 4413
TABLE I: Different positions and orientations collected
  • X, Y and Z are expressed in centimeters. Pitch, roll and yaw orientations are in degrees. Objects refers to the number of objects detected.
@@
A total of 10 different viewpoints were collected. For each viewpoint we release 3 videos of 10 minutes each from different scenes. Given the long time range in which the videos are distributed, we manually selected samples to represent a similar distribution of suburban areas in non-adverse weather and at day-time. Dashcams recorded at an average of 30 frames per second in 720p MP4 format and with a 150° diagonal field of view. We down-sampled the data at 10 fps and extracted the frames undistorted in JPEG format, with a quality factor of 92 and 4:2:0 chroma sub-sampling. This resulted in a total of 360K frames. Note that although the two cameras are simultaneously recording the same scene, no synchronization mechanism was implemented. This can result in the two recordings being shifted by a variable number of frames.
Finally, to foster further research, we also release GPS points and accelerometer data at 1 and 100 Hz, respectively.
IV-A Object Detection
@@
We used a Yolov5 object detector [23], trained on a combination of BDD100K [24] and a private dataset collected with same camera model of this work. We do not expect the object detector to be affected by the viewpoint shifts as both datasets used for training have non-fixed viewpoints. Moreover, from a qualitative analysis, we observe that the combinations in Table I do not induce a significant change in appearance of the objects. This is also confirmed by smaller-scale experiments that showed no evident change in mean average precision between the different viewpoints.
@@
IV-B Homography
As mentioned in Section III, in order to be able to associate a ground truth distance to road objects, we estimated homographies mapping image ground points to a target plane with metric reference system (Fig. 1) that we release alongside the dataset.
IV-B1 Base Camera Homography
@@
IV-B2 Shifted Camera Homography
@@
V Experimental Results
In this section we report the results of our experiments. In particular, we validate our method as ground truth source by comparing it with LIDAR readings on the KITTI dataset in Section V-A. Then, in Section V-B, we use our ground truth method to measure the effects of the different viewpoints on a self-supervised monocular depth estimator.
@@ V Experimental Results
V-A Homography Evaluation on KITTI
The goal of this experiment is to assess the margin of error of our geometry-based strategy by comparing it with LIDAR ground truth. We take the MonoViT model trained on the KITTI dataset [10] for self-supervised monocular depth estimation and measure its absolute relative error (1) using two ground truth sources: laser scans and our strategy.
@@ V-A Homography Evaluation on KITTI
V-B Viewpoint Shift Effects
@@
As mentioned in Section V-A, predictions need to be scaled to metric distances for evaluation. However, in order to highlight the effects of the shifts on the scale perceived by the model, we choose to use the same scaling factor for all positions. This factor is computed exploiting all the objects $O$ in the training set $T$, both in the zero-shot and fine-tuning experiments:
@@
Table II gives the absolute relative errors of the model in the different positions, both in zero-shot and fine-tuned inference. To isolate the effect of the viewpoint, for each position, we test on footage from the same scene recorded by the two cameras simultaneously, and compute the difference in performance between shifted (S) and base (B) camera. The rows with the highest difference between the S and B cameras correspond to the installation positions with the strongest performance degradation.
@@
[] Shifted Dashcam Position Zero-shot Fine-tuned ID X Y Z Pitch Roll Yaw B S S-B B S S-B 0 0 30 0 -4 0.0 0 18.3 18.5 0.2 17.5 17.7 0.3 1 5 0 0 -10 0.0 0 19.3 29.2 9.8 18.0 23.5 5.6 2 5 0 0 16 0.0 0 17.2 25.5 8.3 16.9 21.1 4.2 3 5 0 0 -4 2.5 0 18.3 21.5 3.1 15.1 16.9 1.8 4 5 0 0 -4 5.0 0 21.0 21.2 0.2 16.9 17.6 0.7 5 5 0 0 16 5.0 0 20.4 40.0 19.6 19.7 36.3 16.5 6 -65 -5 -10 5 0.0 -5 19.5 24.1 4.7 16.6 18.7 2.1 7 -65 -5 -10 16 0.0 -5 18.9 46.8 27.9 17.5 44.5 27.0 8 -65 30 -10 5 0.0 -5 18.4 23.6 5.1 17.0 21.7 4.6 9 -65 30 -10 16 0.0 -5 19.0 25.4 6.4 14.8 17.9 3.1
TABLE II: Absolute relative error in different camera positions
  • B and S are the abs-rel error on the base and shifted cameras, respectively. S-B is the degradation computed as difference of errors. The top 2 largest degradations are in bold, the second 2 are underlined.
@@
From the zero-shot results, we observe that the two most affected positions are 7 and 5 (in decreasing order). The combination of pitch and yaw has the most degrading effect, leading to more than doubling the error. Pitch and roll have also the effect of doubling the error of the model. Fine-tuning reduces the error in general, however the degrading effects of the 2 worst positions remain high. In general we observe that pure rotations or translations don’t affect the model as much as a combination of them, with some more impactful than others.
@@
Given the single scale factor used for all positions, we believe that the main effect of viewpoint shifts is the distortion of the scale perceived by the model, which could not have been detected if per-image or per-position scaling was used. In order to better investigate this phenomenon, we compute scaling factors as in (2) on the different positions separately. On the right side of Table II we report the computed scales on the zero-shot experiment where a correlation between performance degradation and perceived scale can be seen. In particular, we observe that higher pitch angles correspond to smaller scales perceived by the model (and therefore a bigger scaling factor necessary to align to GT). In the same way, lower pitch angles generate bigger scales perceived. This observation seems to confirm that indeed depth models use vertical position as depth cue for objects as stated by [9]. In general this suggests that adaptation methods tackling this problem should focus on automatic scaling of models at inference time.
@@
[] Shifted Dashcam Position Zero-shot ID X Y Z Pitch Roll Yaw Error S-B Scale B Scale S Scale S-B 0 0 30 0 -4 0.0 0 0.2 31.3 30.6 -0.6 1 5 0 0 -10 0.0 0 9.8 33.7 26.9 -6.9 2 5 0 0 16 0.0 0 8.3 31.6 43.2 11.6 3 5 0 0 -4 2.5 0 3.1 33.1 31.1 -2.1 4 5 0 0 -4 5.0 0 0.2 36.0 33.6 -2.4 5 5 0 0 16 5.0 0 19.6 33.5 52.5 19.0 6 -65 -5 -10 5 0.0 -5 4.7 33.2 30.4 -2.7 7 -65 -5 -10 16 0.0 -5 27.9 33.2 61.2 28.0 8 -65 30 -10 5 0.0 -5 5.1 31.7 27.3 -4.4 9 -65 30 -10 16 0.0 -5 6.4 34.1 39.9 5.8
TABLE III: Scale perceived in different camera positions
  • Error S-B is the error degradation reported from Table II for reference. Scale B and Scale S are the scales perceived in the base and shifted cameras. Scale S-B is computed as Scale S - Scale B. The top 2 largest degradations are in bold, the second 2 are underlined.
VI Conclusion